BexleyBrooks:

@judethedude I have to retype all this again sorry. The answer is not 32 because that answer is not on the sheet An ice cream company wants to make pints of cookie dough ice cream that sells for $5/pt. The ice cream base is valued at $3.50/pt. How many units of cookie dough valued at $6.25/pt must be mixed with 80 pints of the ice cream base to get the desired cookie dough ice cream?

judethedude:

wow ok I might be dumb.

judethedude:

let me try again

judethedude:

what answers are on the sheet ?

judethedude:

brooks ?

BexleyBrooks:

@judethedude wrote:
what answers are on the sheet ?
Theres 1.2, 20, 10, 40, 100, 96. 8.75, 2.8, 4, and 6

judethedude:

ok thanks

judethedude:

mb I was doin something am I too late ?

Bot1:

it is 32 bro :sob:

euphoriiic:

@bot1 wrote:
it is 32 bro :sob:
they just said it wasn't.

surjithayer:

let x pints valued at 6.25 $/pt be mixed with 80 pints of ice cream valued at $3.50/pt. to make (80+x) pints of ice cream valued at $5/pt. 6.25x+80*3.50=(80+x)*5 6.25 x+280.00=400+5x 6.25x-5x=400-280 1.25x=120 125x=12000 x=12000/125=96 reqd. quantity =96 pt

BexleyBrooks:

@surjithayer wrote:
let x pints valued at 6.25 $/pt be mixed with 80 pints of ice cream valued at $3.50/pt. to make (80+x) pints of ice cream valued at $5/pt. 6.25x+80*3.50=(80+x)*5 6.25 x+280.00=400+5x 6.25x-5x=400-280 1.25x=120 125x=12000 x=12000/125=96 reqd. quantity =96 pt
It's not 32. Sorry that answer is not on the worksheet

judethedude:

its 96 not 32

adrianluvvsyouu2:

\frac{(80 \cdot 3.50) + (x \cdot 6.25)}{80 + x} = 5 Solve this equation and you get 96 pints

adrianluvvsyouu2:

Not sure why LaTeX isn't working, but thats the equation

ihy:

it's 96

nep:

hawk tuah

grimreaper124:

dang

V4MPZSKII:

W